the perimeter of a rectangle is 54m. the length is 6m then twice the width. what is the dimentions
Answer by cleomenius(959)

W = x
L = 2x + 6
2x + 4x + 12 = 54
6x = 42
x = 7 m, this is the width.
14 + 6 = 20 m, this is the length.
2(7) + 2(20) = 54 m.
This does check.
Cleomenius.